ABCDEF plans to borrow EUR 45 million. This company pretends to use the new debt to...

  1. ABCDEF plans to borrow EUR 45 million. This company pretends to use the new debt to repurchase 3,600,000 shares at their market price of EUR 12.50. The yield on the new debt will be 12%. The company has 36 million shares outstanding and EPS of EUR0.60 before the repurchase. The company's tax rate is 30%. What will be the company's EPS after the share repurchase? And if the company ́s tax rate is 35% instead?
Answer #1
Net income before repurchase = 36m*E0.60 = €           21.60 million
Interest on debt = E45m*12% = €             5.40 million
After tax interest = E5.40m*(1-0.35) = €             3.51 million
Net income after repurchase = 21.60-3.51 = €           18.09 million
Number of shares = 36m-3.6m = 32.4 million
EPS after repurchase = 18.09/32.40 = €             0.56